OpenStudy (anonymous):

write an equation in slope-intercept form for the line through the points (-2,4) and (2,0) ??

hero (hero):

1. The slope intercept form of a line is y = mx + b 2. Insert each point, one by one into the formula to get: 4 = -2m + b 0 = 2m + b 3. Add both equations together to get: 4 = 2b 4. Isolate and solve for b to get: 4/2 = b 2 = b 5. Find m: 0 = 2m + b 0 = 2m + 2 -2 = 2m -2/2 = m -1 = m Therefore the slope intercept form of a line is: y = -mx + 2

OpenStudy (anonymous):

thanks :)

hero (hero):

@VeritasVosLiberabit , you should have added 4 to both sides to solve for b

OpenStudy (anonymous):

yeah deleted